# Pushing Cops and Robber on Graphs of Maximum Degree 4

### v1

@@ 3. Graphs with Maximum Degree 4
      Figure 3. An illustrative proof for the Case 1(2)(a) of 3.6 when $|N^{+}(x_{1})|=1$. When $\mathcal{R}$ moves to $x$ in Figure 3(d), it gets trapped.
      First, let $|N^{+}(x_{1})|=1$ and let $N^{+}(x_{1})=\left\{w\right\}$. See Figure 3 for an illustration of the proof. $\mathcal{C}$ begins with pushing $y$. This forces $\mathcal{R}$ to move to $x$, as otherwise, $\mathcal{C}$ will trap $\mathcal{R}$ in the next round by pushing $x$. Next, $\mathcal{C}$ pushes $y_{2}$, forcing to move to $x_{1}$ to avoid getting trapped in the next round. Now, at this point observe that $N^{+}(x_{1})=\left\{w,x_{2},y\right\}$ and $N^{-}(x_{1})=\left\{x\right\}$. Now, $\mathcal{C}$ pushes $x_{1}$ to make $N^{+}(x_{1})=\left\{x\right\}$. Thus, $\mathcal{R}$ is forced to move to $x$ in the next round, as otherwise, it will be trapped by $\mathcal{C}$ at $x_{1}$ by pushing $x$. Hence, it moves to $x$. At this point, observe that $N^{+}(x)=\emptyset$ as both out-neighbors of $x$ (i.e., $x_{1}$ and $x_{2}$) are pushed exactly once and no in-neighbor of $x$ is pushed. Thus, $\mathcal{C}$ gets trapped at $x$.
      Second, let $|N^{+}(x_{1})|=0$. Here, $\mathcal{C}$ begins with pushing $y$, which forces $\mathcal{R}$ to move to $x$. Next, $\mathcal{C}$ pushes $y$ again, returning the graph to its initial configuration. Now, $\mathcal{R}$ can either (i) stay on $x$, (ii) move to $x_{1}$, or (iii) move to $x_{2}$. (i) If $\mathcal{R}$ stays on $x$, then $\mathcal{C}$ pushes $x_{2}$, forcing $\mathcal{R}$ to move to $x_{1}$, and since $|N^{+}(x_{1})|\leqslant 1$ at this point of time, it will be trapped at $x_{1}$ due to Lemma 3.1. (ii) If $\mathcal{R}$ moves to $x_{1}$, then observe that it is trapped since $|N^{+}(x_{1})|=0$. (iii) If $\mathcal{R}$ moves to $x_{2}$, then either $|N^{+}(x_{2})|=1$, i.e., $N^{+}(x_{2})=\left\{x_{1}\right\}$, or $|N^{+}(x_{2})|=2$, i.e., $N^{+}(x_{2})=\left\{x_{1},w^{\prime}\right\}$ for some $w^{\prime}\in V(\overrightarrow{G})$. If $|N^{+}(x_{2})|=1$, then $\mathcal{R}$ gets trapped via an application of Lemma 3.1. If $N^{+}(x_{2})=\left\{x_{1},w^{\prime}\right\}$, then cop pushes $w^{\prime}$, forcing $\mathcal{R}$ to move to $x_{1}$ in the next round. Observe that $|N^{+}(x_{1})|\leqslant 1$ at this point since $|N^{+}(x_{1})|=0$ before $\mathcal{C}$ pushed $w^{\prime}$. Hence, $\mathcal{R}$ will be trapped via an application of Lemma 3.1.
@@
      First, let $N^{+}(x_{1})=\left\{w\right\}$ when $\mathcal{R}$ moved to $u$. $\mathcal{C}$ begins with pushing the vertex $y$, which forces $\mathcal{R}$ to move to $x$ in the next round. Now, observe that $N^{+}(x_{1})=\left\{w,y\right\}$. When $\mathcal{R}$ reaches $x$, $\mathcal{C}$ pushes $x_{2}$, forcing $\mathcal{C}$ to move to $x_{1}$. At this point, observe that $N^{+}(x_{1})=\left\{x_{2},y,w\right\}$ (since the last pushed vertex $x_{2}$ was an in-neighbor of $x_{1}$) and $|N^{+}(x)|=1$ as exactly one of its out-neighbor is pushed but none of its in-neighbor is pushed. When $\mathcal{R}$ reaches $x_{1}$, $\mathcal{C}$ pushes $x_{1}$, which makes $|N^{+}(x_{1})|=1$ (i.e., $N^{+}(x_{1})=\left\{x\right\}$), and hence $\mathcal{R}$ is forced to moved to $x$ in the next round (due to Lemma 3.1) and when $\mathcal{R}$ moves to $x$, $|N^{+}(x)|=1$, and hence it will be trapped at $x$ via an application of Lemma 3.1.
      Second, let $N^{+}(x_{1})=\emptyset$. Again $\mathcal{C}$ begins with pushing $y$, which forces $\mathcal{R}$ to move to $x$ to avoid getting trapped. We push $y$ again returning the graph to its original configuration. Now $\mathcal{R}$ has one of the following three options: (i) $\mathcal{R}$ moves to $x_{1}$, in which case it gets trapped by definition since $|N^{+}(x_{1})|=0$. (ii) $\mathcal{R}$ stays on $x$, in which case $\mathcal{C}$ pushes $x_{2}$ forcing $\mathcal{R}$ to move to $x_{1}$ while $|N^{(}x_{1})|=1$, and hence $\mathcal{R}$ will be trapped due to Lemma 3.1. (iii) $\mathcal{R}$ moves to $x_{2}$. Now, observe that $|N^{+}(x_{2})|\in\left\{1,2,3\right\}$. If $|N^{+}(x_{2})|=1$, then it is trapped using Lemma 3.1. If $|N^{+}(x_{2})|=2$, then $\mathcal{C}$ pushes the out-neighbor of $x_{2}$ other than $x_{1}$, forcing $\mathcal{R}$ to move to $x_{1}$, and at this point, observe that $|N^{+}(x_{1})|\leqslant 1$, and hence $\mathcal{R}$ will be trapped using Lemma 3.1. Finally, if $|N^{+}(x_{2})|=3$, then $\mathcal{C}$ pushes $x_{2}$, forcing $\mathcal{R}$ to move back to $x$, and since $|N^{+}(x)|=1$ (since effectively only $x_{2}$ has been pushed and $x_{2}\in N^{+}(x)$), $\mathcal{R}$ will be trapped using Lemma 3.1
@@
      $x_{1}\neq y_{1}$: First, we show that if $|N^{+}(x_{1})|=0$, then we can trap $\mathcal{R}$. $\mathcal{C}$ begins with pushing $y$, forcing $\mathcal{R}$ to move to $x$. Next, $\mathcal{C}$ pushes $x_{2}$, forcing $\mathcal{R}$ to move to $x_{1}$. At this point, observe that at most one out-neighbor $x_{2}$ of $x_{1}$ is pushed and hence $|N^{+}(x_{1})|=1$, and hence $\mathcal{R}$ will be trapped due to Lemma 3.1. Hence, for the rest of this case, let us assume that $N^{+}(x)=\left\{w_{1}\right\}$. Now, we have three cases depending on whether (i) $|N^{+}(x_{2})|=1$, $|N^{+}(x_{2})|=2$,or (iii) $|N^{+}(x_{2})=3|$.
      (i) First, let $|N^{+}(x_{2})|=1$ and let $N^{+}(x_{2})=\left\{x_{1}\right\}$. Now, $\mathcal{C}$ pushes $y$ forcing $\mathcal{R}$ to move to $x$. Next $\mathcal{C}$ pushes $y$ again, returning the graph to its original configuration. Since both $|N^{+}(x_{1})|=|N^{+}(x_{2})|=1$, if $\mathcal{C}$ moves to any of $x_{1},x_{2}$, it will be trapped due to Lemma 3.1. If it stays on $x$, then $\mathcal{C}$ pushes $x_{1}$, forcing $\mathcal{R}$ to move to $x_{2}$, which is a trap vertex since the only out-neighbor of $x_{2}$ (i.e., $x_{1}$) is pushed exactly once. Hence $\mathcal{R}$ will be trapped in this case.
      (ii)Second, let $|N^{+}(x_{2})|=2$ and let $N+(x_{2})=\left\{x_{1},w_{2}\right\}$ and $N^{-}(x_{2})=\left\{x,w_{3}\right\}$ (possibly $w_{2}$ or $w_{3}$ can be an in-neighbor or out-neighbor of $x_{1}$ or in-neighbor of $x$). Again $\mathcal{C}$ begins with pushing $y$, forcing $\mathcal{R}$ to move to $x$. Now, consider $N^{+}(x_{2})$ at this point of time. If $|N^{+}(x_{2})|=3$ (i.e., $y$ was an in-neighbor of $x_{2}$ before $y$ was pushed), then $\mathcal{C}$ pushes $x_{2}$, forcing $\mathcal{R}$ to move to $x_{1}$. At this point $N^{+}(x_{1})=\left\{w_{1},x_{2}\right\}$. We push $w_{1}$, forcing $\mathcal{R}$ to move to $x_{2}$. If $w_{1}$ was not a neighbor to $x_{2}$, then observe that $|N^{+}(x_{2})|=1$ at this point and hence, $\mathcal{R}$ will be trapped using Lemma 3.1, else, $\mathcal{C}$ pushes $w_{1}$, forcing $\mathcal{R}$ to move to $x$, which has at most one out-neighbor at this point, and hence $\mathcal{R}$ will be trapped due to Lemma 3.1. If $|N^{+}(x_{2})|\leqslant 2$ (after $\mathcal{C}$ has pushed $y$), then $\mathcal{C}$ pushes $x_{1}$ forcing $\mathcal{R}$ to move to $x_{2}$, and since $|N^{+}(x_{2})|\leqslant 1$ after $\mathcal{C}$ pushed $x_{1}$, $\mathcal{R}$ will be trapped at $x_{2}$ due to Lemma 3.1.
      (iii) Third, let $N^{+}(x_{2})=1$. $\mathcal{C}$ begins with pushing $y$ forcing $\mathcal{R}$ to move to $x$. At this point $|N^{+}(x_{2})|\leqslant 2$. Next, $\mathcal{C}$ pushes $x_{1}$, forcing $\mathcal{R}$ to move to $x_{1}$ and ensuring that $|N^{+}(x_{2})|\leqslant 1$ again. Hence, $\mathcal{R}$ will be trapped at $x_{2}$ due to Lemma 3.1.
@@
    $x_{1}\notin\left\{y_{1},y_{2}\right\}$ (i.e., $x_{1}\notin N^{+}(y)$): Here, $\mathcal{C}$ begins with pushing $y$. This forces $\mathcal{R}$ to move to $x$, as otherwise, $\mathcal{C}$ will trap $\mathcal{R}$ by pushing $x$ in the next round. Hence $\mathcal{R}$ moves to $x$. Next, $\mathcal{R}$ pushes $x_{2}$, which forces $\mathcal{R}$ to move to $x_{1}$ in the next round, as otherwise, $\mathcal{C}$ will trap $\mathcal{R}$ in the next round by pushing $x_{1}$. Since $x_{1}$ was a visited vertex, $|N^{+}(x_{1})|\leqslant 1$ before $\mathcal{C}$ started pushing vertices and since none of the pushed vertices, i.e, $y,x_{2}$ are neighbors of $x_{1}$, we have that $|N^{+}(x_{1})|\leqslant 1$ when $\mathcal{R}$ moved to $x_{1}$. Hence, $\mathcal{R}$ will be trapped by $\mathcal{C}$ using Lemma 3.1.
@@
    $x_{1}\in\left\{y_{1},y_{2}\right\}$: Here, since both $x_{1}$ and $y_{1}$ are visited vertices and $x_{1},y_{1}\in N^{+}(y)$, we can assume without loss of generality that $x_{1}=y_{1}$. First, we establish that if $|N^{+}(x_{2})|\neq 2$, then $\mathcal{C}$ can trap $\mathcal{R}$. Since $x\in N^{+}(x_{2})$, if $|N^{+}(x_{2})|\neq 2$, then either $|N^{+}(x_{2})|\leqslant 1$ or $|N^{+}(x_{2})|=3$. In this case (i.e., when $|N^{+}(x_{2})|\neq 2$), $\mathcal{C}$ begins with pushing $y$. This forces $\mathcal{R}$ to move to $x$ as otherwise, $\mathcal{C}$ will trap $\mathcal{R}$ by pushing $x$ in the next round. Next, $\mathcal{C}$ pushes the vertex $y$ again, returning the graph to its initial configuration. Now, $\mathcal{R}$ can either (i) stay at $x$, (ii) move to $x_{1}$, or (iii) move to $x_{2}$. (i) If $\mathcal{R}$ stays on $x$, then $\mathcal{C}$ pushes $x_{2}$, forcing $\mathcal{R}$ to move to $x_{1}$, and since $|N^{+}(x_{1})|\leqslant 1$ at this point of time, it will be trapped at $x_{1}$ due to Lemma 3.1. (ii) If $\mathcal{R}$ moves to $x_{1}$, then $\mathcal{R}$ will be trapped via an application of Lemma 3.1 since $|N^{+}(x_{1})|\leqslant 1$. (iii) If $\mathcal{R}$ moves to $x_{2}$, then either $|N^{+}(x_{2})|\leqslant 1$, or $|N^{+}(x_{2})|=3$. If $|N^{+}(x_{2})|\leqslant 1$, then $\mathcal{R}$ gets trapped via an application of Lemma 3.1. If $|N^{+}(x_{2})|=3$, then $\mathcal{C}$ pushes $x_{2}$ to make $N^{+}(x_{2})=x$, forcing $\mathcal{R}$ to move to $x$ in the next round (as otherwise, $\mathcal{R}$ will be trapped at $x_{2}$ in the next round by pushing $x$). Observe that $|N^{+}(x)|\leqslant 1$ at this point since $|N^{+}(x)|=2$ before $\mathcal{C}$ pushed $x_{2}$ and $x_{2}\in N^{+}(x_{2})$. Hence, $\mathcal{R}$ will be trapped via an application of Lemma 3.1 at $x$. Notice, that, via symmetry, we can also conclude that if $|N^{+}(y_{2})|\neq 2$ (given $x_{1}=y_{1}$), then $\mathcal{C}$ can trap $\mathcal{R}$.
    Therefore, for the rest of the proof of this case, we will assume that $|N^{+}(x_{2})|=2$ and $|N^{+}(y_{2})|=2$. Let $N^{+}(x_{2})=\left\{w_{1},w_{2}\right\}$. Now, we again distinguish the following two cases:
@@
      $y_{2}=x_{2}$: In this case, $N^{+}(x_{2})=\left\{w_{1},w_{2}\right\}$ and $N^{-}(x_{2})=\left\{x,y\right\}$. $\mathcal{C}$ begins with pushing the vertex $y$. This forces $\mathcal{R}$ to move to $x$, as otherwise, $\mathcal{C}$ will trap $\mathcal{R}$ by pushing $x$ in the next round. At this point observe that $N^{+}(x_{2})=\left\{w_{1},w_{2},y\right\}$. Now, $\mathcal{C}$ pushes $x_{1}$, forcing $\mathcal{R}$ to move to $x_{2}$, as otherwise, $\mathcal{C}$ will trap $\mathcal{R}$ by pushing $x_{2}$ in the next round. Now, $\mathcal{C}$ pushes $x_{2}$ ensuring that $N^{+}(x_{2})=\left\{x\right\}$. Hence, in the next round, $\mathcal{R}$ has to move to $x$, as otherwise, $\mathcal{C}$ will trap $\mathcal{R}$ at $x_{2}$ by pushing $x$. Finally, observe that since we have pushed both out-neighnours of $x$ ($x_{1}$ and $x_{2}$) exactly once and have not pushed any in-neighbor of $y$, observe that $\mathcal{R}$ gets trapped at $x$ as $N^{+}(x)=\emptyset$ at this point.
@@
      First, let $\overrightarrow{x_{2}y}\in E(\overrightarrow{G})$ when $\mathcal{R}$ moved to $u$ from $v$. $\mathcal{C}$ begins with pushing $y$, which forces $\mathcal{R}$ to move $x$ in the next round. When $\mathcal{R}$ moves to $x$, observe that since $y$ was an out-neighbor of $x_{2}$ and $|N^{+}(x_{2})|=2$ before we pushed $y$, we have that $|N^{+}(x_{2})|=1$ when $\mathcal{R}$ reaches $x$. Next, $\mathcal{C}$ pushes $x_{1}$, forcing $\mathcal{R}$ to move to $x_{2}$ and since $x_{1}x_{2}\notin E(G)$, we have that $|N^{+}(x_{2})|=1$ when $\mathcal{R}$ moves to $x_{2}$, and hence $\mathcal{R}$ will be trapped at $x_{2}$ due to Lemma 3.1.
      Figure 4. Illustration for the Case 2(2)(b) of 3.6 when $yx_{2}\notin E(G)$.
@@
      Finally, we consider the case when $yx_{2}\notin E(G)$. See Figure 4 for an illustration. Recall that $N^{+}(u)=\left\{x,y\right\}$ and let $N^{-}(u)=\left\{v,v^{\prime}\right\}$. Similarly, let $N^{+}(x)=\left\{x_{1},x_{2}\right\}$, $N^{-}(x)=\left\{u,x^{\prime}\right\}$, $N^{+}(y)=\left\{y_{1},y_{2}\right\},N^{-}(y)=\left\{u,y^{\prime}\right\},N^{+}(x_{2})=\left\{w_{2},w_{3}\right\},N^{-}(x_{2})=\left\{x,w_{1}\right\},N^{+}(x_{1})=\left\{w_{4}\right\},N^{-}(x_{1})=\left\{x,y\right\},N^{+}(y_{2})=\left\{w_{6},w_{7}\right\},N^{-}(y_{2})=\left\{w_{8},y\right\}$.

### v2

@@ 4. Graphs with Maximum Degree 4
      Figure 3. An illustrative proof for the Case 1(2)(a) of ˜4.6 when $|N^{+}(x_{1})|=1$. When $\mathcal{R}$ moves to $x$ in Figure˜3(d), it gets trapped.
      First, suppose $|N^{+}(x_{1})|=1$ and let $N^{+}(x_{1})=\left\{w\right\}$. See Figure 3 for an illustration of the proof. $\mathcal{C}$ begins with pushing $y$, forcing $\mathcal{R}$ to move to $x$. Next, $\mathcal{C}$ pushes $y_{2}$, forcing to move to $x_{1}$. Now, at this point observe that $N^{+}(x_{1})=\left\{w,x_{2},y\right\}$ and $N^{-}(x_{1})=\left\{x\right\}$. Now, $\mathcal{C}$ pushes $x_{1}$ to make $N^{+}(x_{1})=\left\{x\right\}$, forcing $\mathcal{R}$ to move to $x$. At this point, observe that $N^{+}(x)=\emptyset$ as both out-neighbors of $x$ (i.e., $x_{1}$ and $x_{2}$) are pushed exactly once and no in-neighbor of $x$ is pushed. Thus, $\mathcal{R}$ gets trapped at $x$.
      Second, suppose $|N^{+}(x_{1})|=0$. Here, $\mathcal{C}$ begins with pushing $y$, which forces $\mathcal{R}$ to move to $x$. Next, $\mathcal{C}$ pushes $y$ again, returning the graph to its initial configuration. Now, $\mathcal{R}$ can either (i) stay on $x$, (ii) move to $x_{1}$, or (iii) move to $x_{2}$. (i) If $\mathcal{R}$ stays on $x$, then $\mathcal{C}$ pushes $x_{2}$, forcing $\mathcal{R}$ to move to $x_{1}$, and since $|N^{+}(x_{1})|\leqslant 1$ at this point of time, it will be trapped at $x_{1}$ due to Lemma˜4.1. (ii) If $\mathcal{R}$ moves to $x_{1}$, then observe that it is trapped since $|N^{+}(x_{1})|=0$. (iii) If $\mathcal{R}$ moves to $x_{2}$, then either $|N^{+}(x_{2})|=1$, i.e., $N^{+}(x_{2})=\left\{x_{1}\right\}$, or $|N^{+}(x_{2})|=2$, i.e., $N^{+}(x_{2})=\left\{x_{1},w^{\prime}\right\}$ for some $w^{\prime}\in V(\overrightarrow{G})$. If $|N^{+}(x_{2})|=1$, then $\mathcal{R}$ gets trapped via an application of Lemma˜4.1. If $N^{+}(x_{2})=\left\{x_{1},w^{\prime}\right\}$, then cop pushes $w^{\prime}$, forcing $\mathcal{R}$ to move to $x_{1}$. Observe that $|N^{+}(x_{1})|\leqslant 1$ at this point since $|N^{+}(x_{1})|=0$ before $\mathcal{C}$ pushed $w^{\prime}$. Hence, $\mathcal{R}$ will be trapped via an application of Lemma˜4.1.
@@
      First, let $N^{+}(x_{1})=\left\{w\right\}$ when $\mathcal{R}$ moved to $u$. $\mathcal{C}$ begins with pushing the vertex $y$, which forces $\mathcal{R}$ to move to $x$. Now, observe that $N^{+}(x_{1})=\left\{w,y\right\}$. When $\mathcal{R}$ reaches $x$, $\mathcal{C}$ pushes $x_{2}$, forcing $\mathcal{C}$ to move to $x_{1}$. At this point, observe that $N^{+}(x_{1})=\left\{x_{2},y,w\right\}$ (since $x_{1}x_{2}\in E(G)$) and $N^{+}(x)=\left\{x_{1}\right\}$. When $\mathcal{R}$ reaches $x_{1}$, $\mathcal{C}$ pushes $x_{1}$, which makes $N^{+}(x_{1})=\left\{x\right\}$, and hence $\mathcal{R}$ is forced to moved to $x$. When $\mathcal{R}$ moves to $x$, $|N^{+}(x)|=0$, and hence $\mathcal{R}$ is trapped at $x$.
      Second, let $N^{+}(x_{1})=\emptyset$. Again $\mathcal{C}$ begins with pushing $y$, forcing $\mathcal{R}$ to move to $x$. $\mathcal{C}$ pushes $y$ again returning the graph to its original configuration. Now $\mathcal{R}$ has one of the following three options: (i) $\mathcal{R}$ moves to $x_{1}$, in which case it gets trapped by definition since $|N^{+}(x_{1})|=0$. (ii) $\mathcal{R}$ stays on $x$, in which case $\mathcal{C}$ pushes $x_{2}$, forcing $\mathcal{R}$ to move to $x_{1}$ while $|N^{+}(x_{1})|=1$, and hence $\mathcal{R}$ will be trapped due to Lemma˜4.1. (iii) $\mathcal{R}$ moves to $x_{2}$. Now, observe that $|N^{+}(x_{2})|\in\left\{1,2,3\right\}$. If $|N^{+}(x_{2})|=1$, then $\mathcal{R}$ is trapped using Lemma˜4.1. If $|N^{+}(x_{2})|=2$, then $\mathcal{C}$ pushes the out-neighbor of $x_{2}$ other than $x_{1}$, forcing $\mathcal{R}$ to move to $x_{1}$, and at this point, observe that $|N^{+}(x_{1})|\leqslant 1$, and hence $\mathcal{R}$ will be trapped using Lemma˜4.1. Finally, if $|N^{+}(x_{2})|=3$, then $\mathcal{C}$ pushes $x_{2}$, forcing $\mathcal{R}$ to move back to $x$, and since $|N^{+}(x)|=1$ (since effectively only $x_{2}$ has been pushed and $x_{2}\in N^{+}(x)$), $\mathcal{R}$ will be trapped using Lemma˜4.1.
@@
      $x_{1}\neq y_{1}$: First, we show that if $|N^{+}(x_{1})|=0$, then we can trap $\mathcal{R}$. $\mathcal{C}$ begins with pushing $y$, forcing $\mathcal{R}$ to move to $x$. Next, $\mathcal{C}$ pushes $x_{2}$, forcing $\mathcal{R}$ to move to $x_{1}$. At this point, observe that at most one out-neighbor $x_{2}$ of $x_{1}$ is pushed. Thus, $|N^{+}(x_{1})|=1$, and hence $\mathcal{R}$ will be trapped due to Lemma˜4.1. Hence, for the rest of this case, let us assume that $N^{+}(x_{1})=\left\{w_{1}\right\}$. Now, we have three cases depending on whether (i) $|N^{+}(x_{2})|=1$, (ii) $|N^{+}(x_{2})|=2$, or (iii) $|N^{+}(x_{2})|=3$.
      (i) First, suppose $|N^{+}(x_{2})|=1$ and $N^{+}(x_{2})=\left\{x_{1}\right\}$. Now, $\mathcal{C}$ pushes $y$ forcing $\mathcal{R}$ to move to $x$. Next, $\mathcal{C}$ pushes $y$ again, returning the graph to its original configuration. Since both $|N^{+}(x_{1})|=|N^{+}(x_{2})|=1$, if $\mathcal{C}$ moves to any of $x_{1},x_{2}$, it will be trapped due to Lemma˜4.1. If it stays on $x$, then $\mathcal{C}$ pushes $x_{1}$, forcing $\mathcal{R}$ to move to $x_{2}$, which is a trap vertex since the only out-neighbor of $x_{2}$ (i.e., $x_{1}$) is pushed exactly once. Hence $\mathcal{R}$ will be trapped in this case.
      (ii) Second, suppose $|N^{+}(x_{2})|=2$, $N^{+}(x_{2})=\left\{x_{1},w_{2}\right\}$ and $N^{-}(x_{2})=\left\{x,w_{3}\right\}$ (possibly $w_{2}$ or $w_{3}$ can be an in-neighbor or out-neighbor of $x_{1}$ or in-neighbor of $x$). Again $\mathcal{C}$ begins with pushing $y$, forcing $\mathcal{R}$ to move to $x$. Now, consider $N^{+}(x_{2})$ at this point of time. If $|N^{+}(x_{2})|=3$ (i.e., $y=w_{3}$), then $\mathcal{C}$ pushes $x_{2}$, forcing $\mathcal{R}$ to move to $x_{1}$. At this point $N^{+}(x_{1})=\left\{w_{1},x_{2}\right\}$. We push $w_{1}$, forcing $\mathcal{R}$ to move to $x_{2}$. If $w_{1}\neq w_{2}$, then observe that $|N^{+}(x_{2})|=1$ at this point and hence, $\mathcal{R}$ will be trapped using Lemma˜4.1. Else, $\mathcal{C}$ pushes $w_{1}$, forcing $\mathcal{R}$ to move to $x$, which has at most one out-neighbor at this point, and hence $\mathcal{R}$ will be trapped due to Lemma˜4.1. If $|N^{+}(x_{2})|\leqslant 2$ (after $\mathcal{C}$ has pushed $y$), then $\mathcal{C}$ pushes $x_{1}$, forcing $\mathcal{R}$ to move to $x_{2}$, and since $|N^{+}(x_{2})|\leqslant 1$ after $\mathcal{C}$ pushed $x_{1}$, $\mathcal{R}$ will be trapped at $x_{2}$ due to Lemma˜4.1.
      (iii) Third, suppose $N^{+}(x_{2})=1$. $\mathcal{C}$ begins with pushing $y$, forcing $\mathcal{R}$ to move to $x$. At this point $|N^{+}(x_{2})|\leqslant 2$. Next, $\mathcal{C}$ pushes $x_{1}$, forcing $\mathcal{R}$ to move to $x_{2}$ and ensuring that $|N^{+}(x_{2})|\leqslant 1$ again. Hence, $\mathcal{R}$ will be trapped at $x_{2}$ due to Lemma˜4.1.
@@
    $x_{1}\notin\left\{y_{1},y_{2}\right\}$ (i.e., $x_{1}\notin N^{+}(y)$): Here, $\mathcal{C}$ begins with pushing $y$, forcing $\mathcal{R}$ to move to $x$. Next, $\mathcal{R}$ pushes $x_{2}$, forcing $\mathcal{R}$ to move to $x_{1}$. Since $x_{1}$ was a visited vertex, $|N^{+}(x_{1})|\leqslant 1$ before $\mathcal{C}$ started pushing vertices and since none of the pushed vertices, i.e, $y,x_{2}$ are neighbors of $x_{1}$, we have that $|N^{+}(x_{1})|\leqslant 1$ when $\mathcal{R}$ moved to $x_{1}$. Hence, $\mathcal{R}$ will be trapped by $\mathcal{C}$ using Lemma˜4.1.
@@
    $x_{1}\in\left\{y_{1},y_{2}\right\}$: Here, since both $x_{1}$ and $y_{1}$ are visited vertices and $x_{1},y_{1}\in N^{+}(y)$, we can assume without loss of generality that $x_{1}=y_{1}$. First, we establish that if $|N^{+}(x_{2})|\neq 2$, then $\mathcal{C}$ can trap $\mathcal{R}$. Since $x\in N^{-}(x_{2})$, if $|N^{+}(x_{2})|\neq 2$, then either $|N^{+}(x_{2})|\leqslant 1$ or $|N^{+}(x_{2})|=3$. In this case (i.e., when $|N^{+}(x_{2})|\neq 2$), $\mathcal{C}$ begins with pushing $y$, forcing $\mathcal{R}$ to move to $x$. Next, $\mathcal{C}$ pushes the vertex $y$ again, returning the graph to its initial configuration. Now, $\mathcal{R}$ can either (i) stay at $x$, (ii) move to $x_{1}$, or (iii) move to $x_{2}$. (i) If $\mathcal{R}$ stays on $x$, then $\mathcal{C}$ pushes $x_{2}$, forcing $\mathcal{R}$ to move to $x_{1}$, and since $|N^{+}(x_{1})|\leqslant 1$ at this point of time, $\mathcal{R}$ will be trapped at $x_{1}$ due to Lemma˜4.1. (ii) If $\mathcal{R}$ moves to $x_{1}$, then $\mathcal{R}$ will be trapped via an application of Lemma˜4.1 since $|N^{+}(x_{1})|\leqslant 1$. (iii) If $\mathcal{R}$ moves to $x_{2}$, then either $|N^{+}(x_{2})|\leqslant 1$, or $|N^{+}(x_{2})|=3$. If $|N^{+}(x_{2})|\leqslant 1$, then $\mathcal{R}$ gets trapped via an application of Lemma˜4.1. If $|N^{+}(x_{2})|=3$, then $\mathcal{C}$ pushes $x_{2}$ to make $N^{+}(x_{2})=\{x\}$, forcing $\mathcal{R}$ to move to $x$. Observe that $|N^{+}(x)|\leqslant 1$ at this point since, before $\mathcal{C}$ pushed $x_{2}$, we had $|N^{+}(x)|=2$ and $x_{2}\in N^{+}(x)$. Hence, $\mathcal{R}$ will be trapped via an application of Lemma˜4.1 at $x$. Notice, that, via symmetry, we can also conclude that if $|N^{+}(y_{2})|\neq 2$ (given $x_{1}=y_{1}$), then $\mathcal{C}$ can trap $\mathcal{R}$.
    Therefore, for the rest of the proof of this case, we will assume that $|N^{+}(x_{2})|=2$ and $|N^{+}(y_{2})|=2$. Let $N^{+}(x_{2})=\left\{w_{3},w_{2}\right\}$. Now, we again distinguish the following two cases:
@@
      $y_{2}=x_{2}$: In this case, $N^{+}(x_{2})=\left\{w_{3},w_{2}\right\}$ and $N^{-}(x_{2})=\left\{x,y\right\}$. $\mathcal{C}$ begins with pushing the vertex $y$, forcing $\mathcal{R}$ to move to $x$. At this point observe that $N^{+}(x_{2})=\left\{w_{3},w_{2},y\right\}$. Now, $\mathcal{C}$ pushes $x_{3}$, forcing $\mathcal{R}$ to move to $x_{2}$. Now, $\mathcal{C}$ pushes $x_{2}$ ensuring that $N^{+}(x_{2})=\left\{x\right\}$, forcing $\mathcal{R}$ to move to $x$. Finally, observe that since we have pushed both out-neighbors of $x$ ($x_{3}$ and $x_{2}$) exactly once and have not pushed any in-neighbor of $y$, observe that $\mathcal{R}$ gets trapped at $x$ as $N^{+}(x)=\emptyset$ at this point.
@@
      First, let $\overrightarrow{x_{2}y}\in E(\overrightarrow{G})$ when $\mathcal{R}$ moved to $u$ from $v$. $\mathcal{C}$ begins with pushing $y$, forcing $\mathcal{R}$ to move $x$. When $\mathcal{R}$ moves to $x$, observe that since $y$ was an out-neighbor of $x_{2}$ and $|N^{+}(x_{2})|=2$ before we pushed $y$, we have that $|N^{+}(x_{2})|=1$ when $\mathcal{R}$ reaches $x$. Next, $\mathcal{C}$ pushes $x_{1}$, forcing $\mathcal{R}$ to move to $x_{2}$. Since $x_{1}x_{2}\notin E(G)$, we have that $|N^{+}(x_{2})|=1$ when $\mathcal{R}$ moves to $x_{2}$, and hence $\mathcal{R}$ will be trapped at $x_{2}$ due to Lemma˜4.1.
      Figure 4. Illustration for the Case 2(2)(b) of ˜4.6 when $yx_{2}\notin E(G)$.
@@
      Finally, we consider the case when $yx_{2}\notin E(G)$. See Figure˜4 for an illustration. Recall that $N^{+}(u)=\left\{x,y\right\}$ and let $N^{-}(u)=\left\{v,v^{\prime}\right\}$. Similarly, let $N^{+}(x)=\left\{x_{1},x_{2}\right\}$, $N^{-}(x)=\left\{u,x^{\prime}\right\}$, $N^{+}(y)=\left\{y_{1},y_{2}\right\},N^{-}(y)=\left\{u,y^{\prime}\right\},N^{+}(x_{2})=\left\{w_{2},w_{3}\right\},N^{-}(x_{2})=\left\{x,w_{1}\right\},N^{+}(x_{1})=\left\{w_{4}\right\},N^{-}(x_{1})=\left\{x,y,w_{5}\right\},N^{+}(y_{2})=\left\{w_{6},w_{7}\right\},N^{-}(y_{2})=\left\{w_{8},y\right\}$. Here we assumed that $|N^{+}(x_{1})|=1$. If $N^{+}(x_{1})=\emptyset$, then $\mathcal{C}$ can push $y$, forcing $\mathcal{R}$ to move to $x$, and then push $x_{2}$, forcing $\mathcal{R}$ to move to $x_{1}$, at which point Lemma˜4.1 applies.
      Let $U=\left\{v^{\prime},v,u,x^{\prime},x,y^{\prime},y,x_{2},x_{1},y_{2},w_{1},\ldots,w_{8}\right\}$. We note that it is possible that all vertices in $\left\{v^{\prime},v,u,x^{\prime},x,y^{\prime},y,x_{2},x_{1},y_{2},w_{1},\ldots,w_{8}\right\}$ are not distinct. First, we argue that if no vertex of $U$ is pushed and $\mathcal{R}$ moves to either of $x$ or $y$, it will be trapped in at most two cop moves. To see this, let $\mathcal{R}$ move to $x$ (resp. $y$). Then, $\mathcal{C}$ pushes $x_{2}$ (resp. $y_{2}$), which forces $\mathcal{R}$ to move to $x_{1}$ and since $N^{+}(x_{1})=1$, $\mathcal{R}$ will be trapped in the next cop move by pushing $w_{4}$.